OpenStudy (anonymous):

A local hamburger shop sold a combined total of 446 hamburgers and cheeseburgers on Tuesday. There were 54 fewer cheeseburgers sold than hamburgers. How many hamburgers were sold on Tuesday?

OpenStudy (anonymous):

Ok so what you need to do is 446-54 which equals....?

OpenStudy (anonymous):

I did that & it said it was wrong.

OpenStudy (yoloshroom):

Really? Uno momento porfavor

OpenStudy (anonymous):

Hm so that should be the answer though.... 446-56=390 So There were 390 hamburgers sold on Tuesday

OpenStudy (yoloshroom):

careful, u wrote 56 and not 54

OpenStudy (anonymous):

Oh oops :P

OpenStudy (anonymous):

It says that the answer is 250?

OpenStudy (anonymous):

How is that...?

OpenStudy (anonymous):

That doesn't make any sense to me the answer should be 392

OpenStudy (yoloshroom):

Nah it makes sen\se, just have to leave for class now so i cant explain it ;( sorry !

OpenStudy (anonymous):

It's okay...

OpenStudy (anonymous):

x-54 + x = 446 2x-54=446 2x=500 x=250

OpenStudy (anonymous):

Hmmmmmmm

OpenStudy (anonymous):

Is that what it says it is?

OpenStudy (anonymous):

Now that it gave me the explanation, I understand. & yes.

OpenStudy (anonymous):

Okay. Sorry I couldn't help :(